Find the range:
y = 2x + 3 use {0, 1, 2, 3} as the domain

Answer by AnlytcPhil(1810)

Find the range:
y = 2x + 3    use {0, 1, 2, 3} as the domain

Substitute (0) in place of x

y = 2x + 3
y = 2(0) + 3
y = 0 + 3
y = 3

Substitute (1) in place of x

y = 2x + 3
y = 2(1) + 3
y = 2 + 3
y = 5

Substitute (2) in place of x

y = 2x + 3
y = 2(2) + 3
y = 4 + 3
y = 7

Substitute (3) in place of x

y = 2x + 3
y = 2(3) + 3
y = 6 + 3
y = 9


So the range is {3, 5, 7, 9}

Edwin